OpenStudy (blackops2luvr):

The legs of a right triangle are 4 units and 7 units. What is the length of the hypotenuse?

OpenStudy (blackops2luvr):

@iGreen

OpenStudy (blackops2luvr):

@iGreen

OpenStudy (natsent06):

what is 7+4=?

OpenStudy (gorv):

use the pythogorean theorm

OpenStudy (gorv):

hypotenuse =\[\sqrt{base^2+perpendicular^2}\]

OpenStudy (gorv):

\[hypotenuse=\sqrt{4^2+7^2}\]

OpenStudy (gorv):

solve it further ..inside root it is 4^2 and 7^2

OpenStudy (igreen):

\(a^2 + b^2 = c^2\) \(4^2 + 7^2 = c^2\) \(16 + 49 = c^2\) \(65 = c^2\) \(c=\sqrt{65}\) Can you square that? @blackops2luvr

OpenStudy (blackops2luvr):

8.0622577483

OpenStudy (blackops2luvr):

8.0622577483

OpenStudy (blackops2luvr):

oh, the answer is sqrt 65?

OpenStudy (igreen):

Yep.

OpenStudy (blackops2luvr):

The leg of a right triangle is 5 units and the hypotenuse is 8 units. What is the length, in units, of the other leg of the triangle?

OpenStudy (igreen):

That's your answer.

OpenStudy (blackops2luvr):

iGreen ill open it in a new qeustion for ur medal.